# First-order State Space Model for Lightweight Image Super-resolution

### v1

@@ Abstract
State space models (SSMs), particularly Mamba, have shown promise in NLP tasks and are increasingly applied to vision tasks. However, most Mamba-based vision models focus on network architecture and scan paths, with little attention to the SSM module. In order to explore the potential of SSMs, we modified the calculation process of SSM without increasing the number of parameters to improve the performance on lightweight super-resolution tasks. In this paper, we introduce the First-order State Space Model (FSSM) to improve the original Mamba module, enhancing performance by incorporating token correlations. We apply a first-order hold condition in SSMs, derive the new discretized form, and analyzed cumulative error. Extensive experimental results demonstrate that FSSM improves the performance of MambaIR on five benchmark datasets without additionally increasing the number of parameters, and surpasses current lightweight SR methods, achieving state-of-the-art results.
Index Terms:
ODE, state space model, lightweight image super-resolution, error analysis
@@ III-B First-order State Space Model (FSSM)
that maps continuous input $x(t)$ to output $y(t)$ by introducing a state variable $h(t)$, where $x(t),y(t)\in C^{0}(\mathbb{R})$, $B(t),C(t)\in C^{0}(\mathbb{R}^{D})$, and $A(t)\in C^{0}(\mathbb{R}^{D\times D})$. If the parameters of the system do not change over time $t$, it is referred to as a time-invariant SSM. For discrete data, we usually expand the discrete input $\{x_{n}\}$ to continuous $x(t)$, and then solve (1) to get the solution $y(t)$, finally taking time index of $y(t)$ to get discrete output $\{y_{n}\}$.
@@
where $\Delta_{n}=t_{n+1}-t_{n}$. Taking (4) into (2), we can get
$h(t_{n+1})$ $=e^{\Delta_{n}A}h(t_{n})+\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}Bx(s)ds$ (5)
@@
6:   $\overline{B}_{1},\overline{B}_{2}\leftarrow$ discretize($\Delta_{n},B_{n}$)  %use (6)
@@ III-C Approximation
Noticed that the discrete parameter $\overline{B}_{1},\overline{B}_{2}$ has the inverse of $\Delta A$, in order to avoid computing the inverse, we refine the approximation by using the series representation of matrix exponential function (3):
@@
For higher precision, we tried two forms of approximation, and analyzed them with extensive experiments. We called form (7) FSSM and form (8) FSSM+.
@@ III-D Analysis
Since we only have discrete inputs $\{x_{n}\}$ instead of continuous $x(t)$, the missing part contributes to the accumulation of error in the output. Now we analyze the cumulative error of SSM and FSSM, in order to simplify the analysis, we disregard the selective mechanism and assume that ($\Delta,A,B,C$) are time-independent. We consider $\{x_{n}\}$ are sampled from unknown real $x(t)$, $y(t)$ is solved with $x(t)$ by (1), and $\{y_{n}\}$ are calculated from SSM(or FSSM) with $\{x_{n}\}$.
@@
$|y(t_{n})-y_{n}|\leq CL|B|e^{\xi}\Delta^{2}\frac{e^{n\Delta A}-I}{e^{\Delta A}-I}.$ (9)
@@
$|y(t_{n})-y_{n}|\leq\frac{CL|B|e^{\xi}\Delta^{2}}{2}\frac{e^{n\Delta A}-I}{e^{\Delta A}-I}.$ (10)
Noticed that FSSM has the same order of error as SSM, but has a lower upper bound of cumulative error.
@@ IV-C Ablation Study
Increasing the expansion order can reduce the truncation error, and thus reduce the cumulative error. In this section, we compare different expansion order of approximation, FSSM (7) and FSSM+ (8) on benchmark datasets. The PSNR results of x2, x3, x4 light-weight SR presented in Table III show that more expansion terms can improve the performance. Note that with higher expansion, the remainder contains less valid information, so the improvement becomes smaller.

### v2

@@ Abstract
State space models (SSMs), particularly Mamba, have shown promise in NLP tasks and are increasingly applied to vision tasks. However, most Mamba-based vision models focus on network architecture and scan paths, with little attention to the SSM module. In order to explore the potential of SSMs, we modified the calculation process of SSM without increasing the number of parameters to improve the performance on lightweight super-resolution tasks. In this paper, we introduce the First-order State Space Model (FSSM) to improve the original Mamba module, enhancing performance by incorporating token correlations. We apply a first-order hold condition in SSMs, derive the new discretized form, and analyzed cumulative error. Extensive experimental results demonstrate that FSSM improves the performance of MambaIR on five benchmark datasets without additionally increasing the number of parameters, and surpasses current lightweight SR methods, achieving state-of-the-art results. Code is available at https://github.com/Edlinf/FMambaIR.
@@ III-B First-order State Space Model (FSSM)
that maps continuous input $x(t)$ to output $y(t)$ by introducing a state variable $h(t)$, where $x(t),y(t)\in C^{0}(\mathbb{R})$, $B(t),C(t)\in C^{0}(\mathbb{R}^{D})$, and $A(t)\in C^{0}(\mathbb{R}^{D\times D})$. If the parameters of the system do not change over time $t$, it is referred to as a time-invariant SSM. For discrete data, we usually expand the discrete input $\{x_{n}\}$ to continuous $x(t)$, and then solve Eq.(1) to get the solution $y(t)$, finally taking time index of $y(t)$ to get discrete output $\{y_{n}\}$.
@@
where $\Delta_{n}=t_{n+1}-t_{n}$. Taking Eq.(4) into Eq.(2), we can get
$h(t_{n+1})$ $=e^{\Delta_{n}A}h(t_{n})+\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}Bx(s)ds$
$=e^{\Delta_{n}A}h_{n}-A^{-1}\bigg[e^{(t_{n+1}-s)A}Bx(s)\big|_{t_{n}}^{t_{n+1}}$
$\quad-\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}B\frac{1}{\Delta_{n}}(x_{n+1}-x_{n})ds\bigg]$
$=e^{\Delta_{n}A}h_{n}-A^{-1}\bigg[Bx_{n+1}-e^{\Delta_{n}A}Bx_{n}$
$\quad+(\Delta_{n}A)^{-1}(I-e^{\Delta_{n}A})B(x_{n+1}-x_{n})\bigg]$
@@
6:   $\overline{B}_{1},\overline{B}_{2}\leftarrow$ discretize($\Delta_{n},B_{n}$)  %use Eq.(5)
@@ III-C Approximation
Noticed that the discrete parameter $\overline{B}_{1},\overline{B}_{2}$ has the inverse of $\Delta A$, in order to avoid computing the inverse, we refine the approximation by using the series representation of matrix exponential function Eq.(3):
@@
For higher precision, we tried two forms of approximation, and analyzed them with extensive experiments. We called form (6) FSSM and form (7) FSSM+. We also extend the discrete parameter $\overline{B}$ in Mamba as SSM+
$\overline{B}=(1+\frac{1}{2}\Delta_{n}A)\Delta_{n}B_{n}.$ (8)
@@ III-D Analysis
Since we only have discrete inputs $\{x_{n}\}$ instead of continuous $x(t)$, the missing part contributes to the accumulation of error in the output. Now we analyze the cumulative error of SSM and FSSM, in order to simplify the analysis, we disregard the selective mechanism and assume that ($\Delta,A,B,C$) are time-independent. We consider $\{x_{n}\}$ are sampled from unknown real $x(t)$, $y(t)$ is solved with $x(t)$ by Eq.(1), and $\{y_{n}\}$ are calculated from SSM(or FSSM) with $\{x_{n}\}$.
Figure 4: Components of cumulative error.
@@
$|y(t_{n+1})-y_{n}|\leq C\Delta^{2}L\frac{e^{(n+1)\Delta A}-I}{e^{\Delta A}-I}e^{\xi}|B|.$ (9)
@@
$|y(t_{n+1})-y_{n}|\leq\frac{C\Delta^{2}L}{2}\frac{e^{(n+1)\Delta A}-I}{e^{\Delta A}-I}e^{\xi}|B|.$ (10)
See Appendix A for complete proof details. Noticed that FSSM has the same order of error as SSM, but has a lower upper bound of cumulative error.
@@ IV-C Ablation Study
Increasing the expansion order can reduce the truncation error, and thus reduce the cumulative error. In this section, we compare different expansion order of approximation, FSSM (Eq.(6)) and FSSM+ (Eq.(7)) on benchmark datasets. The PSNR results of x2, x3, x4 light-weight SR presented in Table III show that more expansion terms can improve the performance. Note that with higher expansion, the remainder contains less valid information, so the improvement becomes smaller.
@@ V Conclusion
Appendix A Analysis of cumulative error
A-A Mamba
Theorem 1(Cumulative error of Mamba). Consider discrete SSM form (Mamba[2]) with time-independent $(\Delta,A,B,C)$.
$h_{n+1}$ $=e^{\Delta A}h_{n}+\overline{B}x_{n}$ (11)
$y_{n}$ $=Ch_{n+1}$
$h_{0}$ $=0$
$\overline{B}$ $=(\Delta A)^{-1}(e^{\Delta A}-I)\Delta B$
Assume $x(t):\mathbb{R}\rightarrow\mathbb{R}$ is L-Lipschitz and satisfies $x(t_{n})=x_{n}$. There exists $\xi\in(0,\Delta A)$, such that SSM satisfies
$|y(t_{n+1})-y_{n}|\leq C\Delta^{2}L\frac{e^{(n+1)\Delta A}-I}{e^{\Delta A}-I}e^{\xi}|B|.$ (12)
Proof. Let $h(t)$ be the global solution of SSM with boundary condition $h(0)=0$, define $\overline{h}_{n+1}$ as the solution of SSM at $t_{n+1}$ with zero-order hold condition
$h(t_{n+1})$ $=e^{\Delta A}h(t_{n})+\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}Bx(s)ds$
$\overline{h}_{n+1}$ $=e^{\Delta A}h(t_{n})+\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}B\overline{x}(s)ds$
Noticed that $\overline{x}(t)$ holds $x_{n}$ on $t_{n},t_{n+1})$, and $x(t)$ is L-Lipschitz, then
$|x(t)-\overline{x}(t)|\leq\Delta L$
holds on $t_{n},t_{n+1})$. Then
$|h(t_{n+1})-\overline{h}_{n+1}|$ $\leq\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}|B||x(s)-\overline{x}(s)|ds$ (13)
$\leq\Delta L\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}|B|ds$
$=\Delta LA^{-1}(e^{\Delta A}-I)|B|$
$=\Delta^{2}Le^{\xi}|B|\quad\xi\in(0,\Delta A)$
Let $C_{1}=Le^{\xi}|B|$. Noticed that $\overline{h}_{n+1}$ have explicit form
$\overline{h}_{n+1}=e^{\Delta A}h(t_{n})+\overline{B}x_{n}.$
And $h_{n+1}$ is the discrete approximation of $h(t_{n+1})$, with boundary condition $h(t_{n})=h_{n}$ as Eq.(11), then
$|\overline{h}_{n+1}-h_{n+1}|=e^{\Delta A}|h(t_{n})-h_{n}|$ (14)
Then cumulative error at $t_{n+1}$ is
$|h(t_{n+1})-h_{n+1}|$ $\leq|h(t_{n+1})-\overline{h}_{n+1}|+|\overline{h}_{n+1}-h_{n+1}|$ (15)
$\leq e^{\Delta A}|h(t_{n})-h_{n}|+\Delta^{2}C_{1}$
$|h(t_{n+1})-h_{n+1}|+\frac{\Delta^{2}C_{1}}{e^{\Delta A}-I}$ $\leq e^{\Delta A}\left(|h(t_{n})-h_{n}|+\frac{\Delta^{2}C_{1}}{e^{\Delta A}-I}\right)$
$|h(t_{n+1})-h_{n+1}|$ $\leq(e^{(n+1)\Delta A}-I)\frac{\Delta^{2}C_{1}}{e^{\Delta A}-I}.$
Substituting Eq.(15) into Eq.(11) yields Eq.(12).  $\Box$
A-B FMamba
Theorem 2(Cumulative error of FMamba). Consider discrete FSSM form (5) with time-independent $(\Delta,A,B,C)$.
$h_{n+1}$ $=e^{\Delta A}h_{n}+\overline{B}_{1}x_{n}+\overline{B}_{2}x_{n+1}$ (16)
$y_{n}$ $=Ch_{n+1}$
$h_{0}$ $=0$
$\overline{B}_{1}$ $=(\Delta A)^{-2}(\Delta Ae^{\Delta A}-e^{\Delta A}+I)\Delta B$
$\overline{B}_{2}$ $=(\Delta A)^{-2}(e^{\Delta A}-I-\Delta A)\Delta B$
Assume $x(t):\mathbb{R}\rightarrow\mathbb{R}$ is L-Lipschitz and satisfies $x(t_{n})=x_{n}$. There exists $\xi\in(0,\Delta A)$, such that FSSM satisfies
$|y(t_{n+1})-y_{n}|\leq\frac{C\Delta^{2}L}{2}\frac{e^{(n+1)\Delta A}-I}{e^{\Delta A}-I}e^{\xi}|B|.$ (17)
Proof. Let $h(t)$ be the global solution of SSM with boundary condition $h(0)=0$, define $\overline{h}_{n+1}$ as the solution of SSM at $t_{n+1}$ with first-order hold condition
$h(t_{n+1})$ $=e^{\Delta A}h(t_{n})+\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}Bx(s)ds$
$\overline{h}_{n+1}$ $=e^{\Delta A}h(t_{n})+\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}B\overline{x}(s)ds$
The formula of $\overline{x}(t)$ over $t_{n},t_{n+1}]$ is
$\overline{x}(t)=x_{n}+\frac{t-t_{n}}{\Delta_{n}}(x_{n+1}-x_{n})$
Noticed that $x(t)$ is L-Lipschitz, then
$|x(t)-\overline{x}(t)|\leq\Delta L/2$
holds on $t_{n},t_{n+1}]$. Then
$|h(t_{n+1})-\overline{h}_{n+1}|$ $\leq\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}|B||x(s)-\overline{x}(s)|ds$ (18)
$\leq\frac{1}{2}\Delta L\int_{t_{n}}^{t_{n+1}}e^{(t_{n+1}-s)A}|B|ds$
$=\frac{1}{2}\Delta LA^{-1}(e^{\Delta A}-I)|B|$
$=\Delta^{2}Le^{\xi}|B|/2\quad\xi\in(0,\Delta A)$
Let $C_{1}=Le^{\xi}|B|/2$. Noticed that $\overline{h}_{n+1}$ have explicit form
$\overline{h}_{n+1}=e^{\Delta A}h(t_{n})+\overline{B}_{1}x_{n}+\overline{B}_{2}x_{n+1}.$
And $h_{n+1}$ is the discrete approximation of $h(t_{n+1})$, with boundary condition $h(t_{n})=h_{n}$, as Eq.(16), then
$|\overline{h}_{n+1}-h_{n+1}|=e^{\Delta A}|h(t_{n})-h_{n}|$ (19)
Then cumulative error at $t_{n+1}$ is
$|h(t_{n+1})-h_{n+1}|$ $\leq|h(t_{n+1})-\overline{h}_{n+1}|+|\overline{h}_{n+1}-h_{n+1}|$ (20)
$\leq e^{\Delta A}|h(t_{n})-h_{n}|+\Delta^{2}C_{1}$
$|h(t_{n+1})-h_{n+1}|+\frac{\Delta^{2}C_{1}}{e^{\Delta A}-I}$ $\leq e^{\Delta A}\left(|h(t_{n})-h_{n}|+\frac{\Delta^{2}C_{1}}{e^{\Delta A}-I}\right)$
$|h(t_{n+1})-h_{n+1}|$ $\leq(e^{(n+1)\Delta A}-I)\frac{\Delta^{2}C_{1}}{e^{\Delta A}-I}.$
Substituting Eq.(20) into Eq.(16) yields Eq.(17).  $\Box$